Recent Maintenance – NIL

NIL operational status:

2022.04.28

Telemetry was down day 2022,118 (4/28) to 2022,131 (5/11). Station operator restored link to the station.

2021.09.12

The station operator replaced the NIL-10 Q330 on 2021,255 (9/12).

2021.06.11

The station operator replaced the NIL-00 Q330 on 2021,162 (June 11).  However, all channels from the KS54000 are nonseismic bit noise. IDA engineers are working with the station operator to resolve this issue.

2021.05.19

The station is back online after the station operator fixed and connected the GPS cable to the NIL10 Q330.

2021.05.10

The station is down as of day 2021,130 due to communication issues between the NIL00 Q330 and the station computers and GPS issues.  IDA engineers have sent instructions to examine the Q330 and connect the spare GPS cable to NIL10 Q330.  All channels from the NIL00 Q330 are out until the Q330 digitizer is replaced.

2020.07.03

The station is back online after the station operator visited the station and replaced the EdgeRouter on day 2020,185 (July 3).

2020.05.22

The station is unreachable since May 9.  Failure of a router at the station is suspected and IDA engineers are working with the station operator to resolve this issue.

2019.02.06

The station operator reports replacing the IPS (integrated power system) and nil2 station computer.  An IDA engineer sent additional instructions to install a new EdgeRouter and Advantech switch.

2017.04.21

The station continues to have power problems.  The station operator reports faulty IPS (integrated power system) on 2017.04.18.   IDA engineers are planning shipment of power equipment to resolve this issue.

2016.10.21

Power to the station has become more stable, resulting in less data gapping. IDA engineers are still working with the station operator to resolve the power problems.

2016.09.12

The station is down as of day 2016,254-02:09 due to power problems at the station.  IDA engineers are working with the station operator to resolve the issue.

2016.07.15

The GPS clocks are locked again as of day 2016,196-23:49.

2016.02.24

The GPS clocks are unlocked as of day 2016,054-08:17.

2013.08.20

All channels from the KS-54000 were restored as of day
2013,234-06:15 after the station operator repaired the SIB.

2013.08.15

All channels from the KS-540000 seismometer are non-seismic as of day
2013,227-03:45 due to problems with the seismometer interface box (SIB).
IDA engineers are working with the station operator to repair or replace the SIB.
2013.04.23

There are consistent pulses on the horizontal channels from the KS-54000. The Trillium seismometer has also been rotated from it’s original position. The station operator reports heavy construction in the area that may be causing the channel pulses/noise.

2013.04.17

All channels from the KS-54000 seismometer were restored after the station operator repaired the seismometer interface box. Data are seismic as of day 2013,107-10:20.

2013.03.14

All channels from the Trillium 240 are seismic after the masses for the sensor were re-centered on day 2013,073-18:38. KS-54000 channels are still non-seismic.

2013.03.07

The station is down due to major damage to station equipment as a result of flooding in the vault.

2013.03.06

There are problems with communication from the digitizers to the station computer. There is no data received since day 2013,065-06:17.

2013.03.05

Telemetry was restored after the station operator replaced the modem. All channels from the KS-54000 seismometer are non-seismic since day 2013,059-01:22 for reasons yet to be determined.

2013.03.04

Telemetry to the station is down since day 2013,058-00:04. The station operator reports problems with the modem. IDA engineers are working with the station operator to resolve this problem.

2012.11.26

Telemetry to the station was restored after the internet connection was fixed.

2012.11.13

Telemetry to the station remains down. However, the station operator reports the station is still running, and will send any missing data on a thumb drive.

2012.11.05

Telemetry to the station is down as of day 2012,299-08:20. The station operator reports an internet problem that they will solve in the next couple days.

2012.09.26

The station operator reports replacing the NIL-10 DAS with the one shipped from UCSD. The circuit breakers at the station continue to trip for reasons yet to be determined.

2012.08.15

IDA station NIL was down from day 2012,223-18:28 to 2012,228-04:19 due to a power outage caused by lightning.

2012.07.13

The NIL-10 DAS is causing data gaps since day 2012,195-09:00. A replacement DAS is prepared to ship to the station pending an export license.

2012.03.30

IDA station NIL has been upgraded to a Quanterra Q330HR system. Seismometers include KS54000 and Trillium 240 seismometers, and a strong motion sensor.

2009.05.02

The station is down as of about day 2009,123-10:26 due to damage to the DAS and ARS caused by lightning.

2009.01.06

The high gain short period channels from the Trillium seismometer were saturated from about day 2009,004-22:50 to 2009,006-18:00.

2008.11.07

The station was down from about day 2008,308 to 2008,312 due to damage to a cable between the ARS and the DAS caused by construction near the station.

2008.10.09

The station is back online as of about day 2008,282-06:23 after the short period channels from the Trillium seismometer came back into normal operating range.

2008.09.26

The station is down. The short period channels from the Trillium seismometer are saturated as of about day 2008,263. An IDA engineer will give the station operator directions to adjust the sensor.

2008.07.07

The station is back on-line as of about day 2008,186 after the station operator replaced the ISP.

2008.07.01

The station is down as of about day 2008,179 due to problems with the ISP.

2008.05.26

The station is back online after the station operator installed a new Trillium seismometer shipped from UCSD.

2006.12.10

The station is down as of about 344-04:55 due to a thunderstorm that damaged equipment including the DAS, ARS, and ISP.

2005.12.21

The station is back up as of day 355 (Dec.21).

2005.12.16

The station is down for reasons yet to be determined.

2005.12.05

The station is back up and data acquisition resumed.

2005.12.03

The station is down as of day 335 (Dec.1) due to communication problems between the DAS and the ISP.

2005.03.07

The station is up since day 033. The vertical and E-W channels for the KS-54000 seismometer are out.

2004.12.20

The station is down due to problems with the DAS that are yet to be determined.

2004.06.08

Station operator reported that DAT tape drive failed to record data.

2004.06.01

The one remaining 24-bit channel has been acquring data from the N-S sensor but was wrongly assigned to BBE. This problem occured roughly around 2004044 and has been corrected during processing stage. Telemetry broadband data between 2004044 to 2004153 stay unchanged, ie they are assigned to incorrected channels BBE.

2004.03.08

Station is back up. Station operator requested to reboot the DAS after restoring proper date/time for the station processor.

2004.02.17

Station is down as of 2004:040-10:24 (Feb.9) due to problems with the DAS.

2003.05.25

Station operator notes failed RT303 card for the vertical channels.

2002.09.06

New EPROMs installed to correct DSP problem.

2002.04.22

The channels from the KS-54000 vertical are non-seismic. Tape data are currently not available.